Seether Co. wants to issue new 15-year bonds for some much-needed expansion projects. The company currently has 11.6 percent coupon bonds on the market that sell for $1,423.02, make semiannual payments, and mature in 15 years. What coupon rate should the company set on its new bonds if it wants them to sell at par?
Answer #1

Price = 1423.02
Coupon rate = 11.6%
Coupon = Coupon Rate * Par Value /2 = 11.6%*1000/2 = 58

Maturity = 15 years

Number of Periods = 15*2 = 30
Par value = 1000

YTM using Financial calculator

N = 30; PMT = 58; FV = 1000; PV = -1423.02; CPT I/Y

Rate = 3.5%

YTM = 2*3.5% = 7%

So coupon rate has to be 7% for price to be at par.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well